Use the distributive property to rewrite and solve this problem 6(399)
Found 2 solutions by stanbon, MathTherapy:
Answer by stanbon(75887)

Use the distributive property to rewrite and solve this problem 6(399)
---------------------------
6(500-1) = 3000-6 = 2994
===============================
Cheers,
Stan H.
Reply to stanbon@comcast.net

Answer by MathTherapy(10858)


6(300 + 99)
1,800 + 594 = 2,394